If P= x 3 - 1/ x 3 and Q=x-1/x , x belongs from 0 to infinity , find the minimum value of P/Q 2

If P=x3 - 1/x3 and Q=x-1/x , x belongs from 0 to infinity , find the minimum value of P/Q2

Grade:Select Grade

6 Answers

udit maniyar
8 years ago
P = x- 1/x3 =(x-1/x)(x+1/x2 +1)
Q=x – 1/x 
 
 
P/Q  when simplified it comes in the form of (x-1/x)+3/(x-1/x) 
then we can apply A.M greater than or equal to G.M
we get the minimum value of p/q2 as 2(3)1/2
 
udit maniyar
8 years ago
if you feel the answer is right then approve it
grenade
8 years ago
but i was getting answer 2
udit maniyar
8 years ago
did u get this after simplification p/q2 =  (x –  1/x) + 3/(x – 1/x) 
 
grenade
8 years ago
askiitians faculty pl. help us 
Tulika Sureka
8 years ago
2(3)1/2 is the correct